MID-IR LASER SPECTROSCOPY OF COF_{2} BETWEEN 1953.5 CM^{−1} AND 1956.0 CM^{−1}

\underline{PATRICK J. McCANN}, I-NA CHAO, and KHOSROW NAMJOU, School of Electrical and Computer Engineering, 202 West Boyd Street, Room 219, University of Oklahoma, Norman, OK 73019.

This talk will describe the recent development of a user-friendly mid-IR laser spectrometer, one that does not need liquid nitrogen refills, and its use in the measurement of carbonyl fluoride (COF_{2}). COF_{2} is a reaction product that is formed during fluorocarbon plasma etching of oxide insulators. Its measurement can provide etching endpoint detection information for high yield semiconductor manufacturing. A IV-VI semiconductor mid-IR laser with single mode emission in the 5.2 micron spectral range was selected to probe the R-branch of the ν_{1} band of COF_{2}. Individual rovibrational lines for COF_{2} as well as an H_{2}O line at 1955.0 cm^{−1} were easily resolved with a spectral resolution of better than 0.005 cm^{−1} using second harmonic detection. Gas samples containing COF_{2} were generated by burning Teflon (PTFE) with a propone torch and collected into a 10 cm long gas cell equipped with CaF_{2} windows held at a pressure of 200 Torr. COF_{2} concentration values higher than 200 ppm were measured using this setup, which had a minimum detection sensitivity of about 50 ppm. In addition to obtaining laser absorption spectra between 1953.5 cm^{−1} and 1956.0 cm^{−1}, concentration values for both COF_{2} and H_{2}O were obtained every 100 msec. Exponential fits to the data showed that both COF_{2} and H_{2}O molecules had lifetimes of about 12 minutes suggesting that they react with each other to form HF and CO_{2}.